OpenStudy (anonymous):

32/48 1/3 7/12 into same demoniator and then order from least to greatest

OpenStudy (lalaly):

\[\frac{32}{48}\]\[\frac{1}{3} \times \frac{16}{16} = \frac{16}{48}\]\[\frac{7}{12} \times \frac{4}{4} = \frac{28}{48}\]

OpenStudy (lalaly):

from least to greatest \[\frac{1}{3} < \frac{7}{12} < \frac{32}{48}\]